# Missing Data Imputation by Reducing Mutual Information with Rectified Flows

### v2

@@ Appendix B Proof of Theorem 1
First, we introduce the marginal preserving property for the imputation ODE.
@@
$\mathbf{Z}_{0}\stackrel{{\scriptstyle d}}{{=}}\mathbf{X}_{0,1-\mathbf{m}}\mid\left\{\mathbf{X}_{0,\mathbf{m}}=\mathbf{x}_{\mathbf{m}},\mathbf{M}_{0}=\mathbf{m}\right\}$
@@
See the Appendix C for the proof.
Now we derive the main result. Lemma 1 shows that under the initial condition,
$\mathbf{Z}_{1}\stackrel{{\scriptstyle d}}{{=}}\mathbf{X}_{1,1-\mathbf{m}}|\mathbf{X}_{1,\mathbf{m}}=\mathbf{x}_{\mathbf{m}}.$ (12)
$\mathbf{Z}_{1}$ is our imputer $\mathbf{g}_{t}$ at iteration $t$. According to (7), the imputer is constructed as $\mathbf{g}_{t}(\cdot;\mathbf{X}^{(t-1)}_{\mathbf{m}},\mathbf{M})$, thus
$\mathbf{Z}_{1}\stackrel{{\scriptstyle d}}{{=}}\mathbf{X}_{\mathbf{g}^{*},\mathbf{m}}^{(t)}\mid\left\{\mathbf{X}_{1-\mathbf{m}}^{(t-1)}=\mathbf{x}_{1-\mathbf{m}},\mathbf{M}_{0}=\mathbf{m}\right\}.$ (13)
Combining (13) and (12), we obtain the following equality
$\mathbb{P}_{\mathbf{X}^{(t)}_{\mathbf{g}^{*},1-\mathbf{m}}\mid\mathbf{X}^{(t-1)}_{\mathbf{m}}=\mathbf{x}_{\mathbf{m}},\mathbf{M}=\mathbf{m}}=\mathbb{P}_{\mathbf{X}^{(t-1)}_{1-\mathbf{m}}\mid\mathbf{X}^{(t-1)}_{\mathbf{m}}=\mathbf{x}_{\mathbf{m}}}.$
Our imputation ODE does not change components where $m_{j}=1$, so $\mathbf{X}^{(t-1)}_{\mathbf{m}}=\mathbf{X}^{(t)}_{\mathbf{m}}$. Thus,
$\mathbb{P}_{\mathbf{X}^{(t)}_{\mathbf{g}^{*},1-\mathbf{m}}\mid\mathbf{X}^{(t-1)}_{\mathbf{m}}=\mathbf{x}_{\mathbf{m}},\mathbf{M}=\mathbf{m}}=\mathbb{P}_{\mathbf{X}^{(t)}_{\mathbf{g}^{*},1-\mathbf{m}}\mid\mathbf{X}^{(t)}_{\mathbf{m}}=\mathbf{x}_{\mathbf{m}},\mathbf{M}=\mathbf{m}}.$
The above two equalities imply that $p_{\mathbf{g}^{*}}(\cdot|\mathbf{x}_{\mathbf{m}},\mathbf{m})=q(\cdot|\mathbf{x}_{\mathbf{m}}).$ According to Proposition 2, we have shown that $\mathbf{g}^{*}$ is an optimal imputer that minimizes $\mathrm{D}\left[\mathbb{P}_{\mathbf{X}^{(t)}(\mathbf{g}),\mathbf{M}}|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right]$. ∎
Appendix C Proof of Lemma 1
@@
$\Xi(\mathbf{u},\mathbf{w},\mathbf{m})=\{\mathbf{X}_{0,\mathbf{m}}=\mathbf{u},\mathbf{X}_{1,\mathbf{m}}=\mathbf{w},\mathbf{M}_{0}=\mathbf{m}\}.$
@@
$\mathbf{v}_{\mathbf{m}}^{*}(\mathbf{y},\mathbf{u},\mathbf{w},\mathbf{m},\tau):=\mathbb{E}\Bigl{[}\dot{\mathbf{X}}_{\tau,1-\mathbf{m}}\,\Big{|}\,\mathbf{X}_{\tau,1-\mathbf{m}}=\mathbf{y},\Xi(\mathbf{u},\mathbf{w},\mathbf{m})\Bigr{]}.$
We denote the density of random variable $\mathbf{X}_{\tau,1-\mathbf{m}}\mid\Xi(\mathbf{u},\mathbf{w},\mathbf{m})$ as $p_{\tau}(\mathbf{y}\mid\mathbf{u},\mathbf{w},\mathbf{m}).$
@@
We can see that due to the independence $\mathbf{X}_{1}\perp\!\!\!\!\perp\mathbf{X}_{0},\mathbf{M}_{0}$,
$\mathbf{X}_{0,1-\mathbf{m}}\mid\mathbf{X}_{0,\mathbf{m}},\mathbf{X}_{1,\mathbf{m}},\mathbf{M}_{0}=\mathbf{X}_{0,1-\mathbf{m}}\mid\mathbf{X}_{0,\mathbf{m}},\mathbf{M}_{0},$
@@
$\mathbf{X}_{1,1-\mathbf{m}}\mid\mathbf{X}_{0,\mathbf{m}},\mathbf{X}_{1,\mathbf{m}},\mathbf{M}_{0}=\mathbf{X}_{1,1-\mathbf{m}}\mid\mathbf{X}_{1,\mathbf{m}},$
@@
The density function $p_{\tau}(\mathbf{y}\mid\mathbf{u},\mathbf{w},\mathbf{m})$ satisfies the continuity equation
$\partial_{\tau}p_{\tau}(\mathbf{y}\mid\mathbf{u},\mathbf{w},\mathbf{m})+\nabla_{\mathbf{y}}\cdot\Big{(}p_{\tau}(\mathbf{y}\mid\mathbf{u},\mathbf{w},\mathbf{m})\,\mathbf{v}^{*}(\mathbf{y},\mathbf{u},\mathbf{w},\mathbf{m},\tau)\Big{)}=0$ (14)
@@
$\mathbb{E}\big{[}h(\mathbf{X}_{\tau,1-\mathbf{m}})\mid\Xi(\mathbf{u},\mathbf{w},\mathbf{m})\big{]}=\int_{\mathbb{R}^{d_{\mathbf{m}}}}h(\mathbf{y})\,p_{\tau}(\mathbf{y}\mid\mathbf{u},\mathbf{w},\mathbf{m})\,\mathrm{d}\mathbf{y}$ (15)
@@
$\partial_{\tau}\mathbb{E}\big{[}h(\mathbf{X}_{\tau,1-\mathbf{m}})\mid\Xi(\mathbf{u},\mathbf{w},\mathbf{m})\big{]}=\int h(\mathbf{y})\,\partial_{\tau}p_{\tau}(\mathbf{y}\mid\mathbf{u},\mathbf{w},\mathbf{m})\,\mathrm{d}\mathbf{y}.$ (16)
On the other hand, we can see that $\mathbf{X}_{\tau}$ is a function of $\mathbf{X}_{0}$ and $\mathbf{X}_{1}$, thus using the law of unconscious statistician:
$\partial_{\tau}\mathbb{E}\big{[}h(\mathbf{X}_{\tau,1-\mathbf{m}})\mid\Xi(\mathbf{u},\mathbf{w},\mathbf{m})\big{]}$ $=\partial_{\tau}\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{1}}\big{[}h(\mathbf{X}_{\tau,1-\mathbf{m}})\mid\Xi(\mathbf{u},\mathbf{w},\mathbf{m})\big{]}$
$=\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{1}}\big{[}\partial_{\tau}h(\mathbf{X}_{\tau,1-\mathbf{m}})\mid\Xi(\mathbf{u},\mathbf{w},\mathbf{m})\big{]}=\mathbb{E}\big{[}\partial_{\tau}h(\mathbf{X}_{\tau,1-\mathbf{m}})\mid\Xi(\mathbf{u},\mathbf{w},\mathbf{m})\big{]}.$
@@
$\partial_{\tau}\mathbb{E}\big{[}h(\mathbf{X}_{\tau,1-\mathbf{m}})\mid\Xi(\mathbf{u},\mathbf{w},\mathbf{m})\big{]}=$ $\mathbb{E}\big{[}\partial_{\tau}h(\mathbf{X}_{\tau,1-\mathbf{m}})\mid\Xi(\mathbf{u},\mathbf{w},\mathbf{m})\big{]}$ (17)
@@
Now we equate the two representations ((16) and (22)) for $\partial_{\tau}\mathbb{E}[h(\mathbf{X}_{\tau,1-\mathbf{m}})\mid\Xi(\mathbf{u},\mathbf{w},\mathbf{m})]$:
@@
Since this equality holds for every smooth, compactly supported test function $h(\mathbf{y})$, the fundamental lemma of the calculus of variations implies that, in the sense of distributions,
@@
We can see that with the initial condition $\mathbf{Z}_{0}=\mathbf{X}_{0,1-\mathbf{m}}\mid\Xi(\mathbf{u},\mathbf{w},\mathbf{m})$, Lemma 3 says $\mathbf{Z}_{1}=\mathbf{X}_{1,1-\mathbf{m}}\mid\Xi(\mathbf{u},\mathbf{w},\mathbf{m})$. From Lemma 2, we can see that $\mathbf{Z}_{0}=\mathbf{X}_{0,1-\mathbf{m}}\mid\mathbf{X}_{0,\mathbf{m}}=\mathbf{u},\mathbf{M}_{0}=\mathbf{m}$ and $\mathbf{Z}_{1}=\mathbf{X}_{1,1-\mathbf{m}}\mid\mathbf{X}_{1,\mathbf{m}}=\mathbf{w}$. Letting $\mathbf{u}=\mathbf{w}=\mathbf{x}_{\mathbf{m}}$, we obtain the desired results in Lemma 1.
Appendix D Proof of Proposition 2
@@
Since the sequential imputer (2) never changes $\mathbf{X}_{\mathbf{m}}^{(t-1)}$ nor $\mathbf{M}$, we have
@@
where the marginal density $p(\mathbf{x}_{1-\mathbf{m}},\mathbf{m})$ does not depend on $\mathbf{g}$. Now we show that the KL divergence only depends on $\mathbf{g}$ through the KL divergence $\mathrm{D}[p_{\mathbf{g}}(\cdot|\mathbf{x}_{1-\mathbf{m}},\mathbf{m})|q(\cdot|\mathbf{x}_{1-\mathbf{m}})]$.
$\mathrm{D}\left[\mathbb{P}_{\mathbf{X}^{(t)}(\mathbf{g}),\mathbf{M}}|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right]$
$=\mathbb{E}_{(\mathbf{x},\mathbf{m})\sim(\mathbf{X}_{\mathbf{g}}^{(t)},\mathbf{M})}\left[\log\frac{p_{\mathbf{g}}(\mathbf{x},\mathbf{m})}{q(\mathbf{x})p(\mathbf{m})}\right]$
$=\mathbb{E}_{(\mathbf{x},\mathbf{m})\sim(\mathbf{X}_{\mathbf{g}}^{(t)},\mathbf{M})}\left[\log\frac{p_{\mathbf{g}}(\mathbf{x}_{1-\mathbf{m}},\mathbf{x}_{\mathbf{m}},\mathbf{m})}{q(\mathbf{x}_{1-\mathbf{m}},\mathbf{x}_{\mathbf{m}})p(\mathbf{m})}\right]$
$=\mathbb{E}_{(\mathbf{x},\mathbf{m})\sim(\mathbf{X}_{\mathbf{g}}^{(t)},\mathbf{M})}\left[\log\frac{p_{\mathbf{g}}(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}},\mathbf{m})}{q(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}},\mathbf{m})}\right]+\mathbb{E}_{(\mathbf{x},\mathbf{m})\sim(\mathbf{X}_{\mathbf{g}}^{(t)},\mathbf{M})}\left[\log\frac{p(\mathbf{x}_{\mathbf{m}},\mathbf{m})}{q(\mathbf{x}_{\mathbf{m}},\mathbf{m})p(\mathbf{m})}\right]$
$=\sum_{\mathbf{m}\in\{0,1\}^{d}}p(\mathbf{m})\mathbb{E}_{\mathbf{x}\sim\mathbf{X}_{\mathbf{g}}^{(t)}|\mathbf{M}=\mathbf{m}}\left[\log\frac{p_{\mathbf{g}}(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}},\mathbf{m})}{q(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}})}\right]+\mathrm{const.}$
$=\sum_{\mathbf{m}\in\{0,1\}^{d}}p(\mathbf{m})\underbrace{\int p(\mathbf{x}_{\mathbf{m}}|\mathbf{m})\mathrm{D}[p_{\mathbf{g}}(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}},\mathbf{m})|q(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}})]\mathrm{d}\mathbf{x}_{\mathbf{m}}}_{A_{\mathbf{g}}}+\mathrm{const.},$

### v3

@@ Appendix C Proof of Theorem 1
First, we introduce the following lemma, based on the marginal preserving property of rectified flow.
@@
$\mathbf{Z}_{0}\stackrel{{\scriptstyle d}}{{=}}\mathbf{X}_{0,1-\mathbf{m}}|\left\{\mathbf{X}_{0,\mathbf{m}}=\mathbf{x}_{\mathbf{m}},\mathbf{M}_{0}=\mathbf{m}\right\}$ (21)
@@
See the Appendix D for the proof.
Now we derive the main result. Lemma 1 shows that under the initial condition, the solution $\mathbf{Z}_{1}$ is
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{Z}_{1}}$ $\stackrel{{\scriptstyle d}}{{=}}\mathbf{X}_{1,1-\mathbf{m}}|\left\{\mathbf{X}_{1,\mathbf{m}}=\mathbf{x}_{\mathbf{m}}\right\}$
$=\mathbf{X}_{1-\mathbf{m}}^{(t-1)}|\left\{\mathbf{X}_{\mathbf{m}}^{(t-1)}=\mathbf{x}_{\mathbf{m}}\right\}$ (22)
Moreover, $\mathbf{Z}_{1}$ is also our imputer $\mathbf{g}_{t}$, so by construction (9),
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\mathbf{Z}_{1}}\stackrel{{\scriptstyle d}}{{=}}\mathbf{X}_{\mathbf{g}^{*},\mathbf{m}}^{(t)}|\left\{\mathbf{X}_{1-\mathbf{m}}^{(t-1)}=\mathbf{x}_{1-\mathbf{m}},\mathbf{M}_{0}=\mathbf{m}\right\}.$ (23)
Combining (23) and (C), we obtain the following equality
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}{\mathbb{P}_{\mathbf{X}^{(t)}_{\mathbf{g}^{*},1-\mathbf{m}}|\mathbf{X}^{(t-1)}_{\mathbf{m}}=\mathbf{x}_{\mathbf{m}},\mathbf{M}=\mathbf{m}}}}=\mathbb{P}_{\mathbf{X}^{(t-1)}_{1-\mathbf{m}}|\mathbf{X}^{(t-1)}_{\mathbf{m}}=\mathbf{x}_{\mathbf{m}}}.$ (24)
The imputation ODE does not change observed components where $m_{j}=1$, so $\mathbf{X}^{(t-1)}_{\mathbf{m}}=\mathbf{X}^{(t)}_{\mathbf{m}}$. We have
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathbb{P}_{\mathbf{X}^{(t)}_{\mathbf{g}^{*},1-\mathbf{m}}|\mathbf{X}^{(t-1)}_{\mathbf{m}}=\mathbf{x}_{\mathbf{m}},\mathbf{M}=\mathbf{m}}}=\mathbb{P}_{\mathbf{X}^{(t)}_{\mathbf{g}^{*},1-\mathbf{m}}|\mathbf{X}^{(t)}_{\mathbf{m}}=\mathbf{x}_{\mathbf{m}},\mathbf{M}=\mathbf{m}}.$ (25)
(24) and (25) imply that $p_{\mathbf{g}^{*}}(\cdot|\mathbf{x}_{\mathbf{m}},\mathbf{m})=q(\cdot|\mathbf{x}_{\mathbf{m}}).$ According to Proposition 2, we have shown that $\mathbf{g}^{*}$ is an optimal imputer that minimizes $\mathrm{D}\left[\mathbb{P}_{\mathbf{X}^{(t)}(\mathbf{g}),\mathbf{M}}\|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right]$. ∎
Appendix D Proof of Lemma 1
@@
$\Xi(\mathbf{u},\mathbf{w},\mathbf{m})=\{\mathbf{X}_{0,\mathbf{m}}=\mathbf{u},\mathbf{X}_{1,\mathbf{m}}=\mathbf{w},\mathbf{M}_{0}=\mathbf{m}\}.$ (26)
@@
$\mathbf{v}_{\mathbf{m}}^{*}(\mathbf{y},\mathbf{u},\mathbf{w},\mathbf{m},\tau):=\mathbb{E}\Bigl[\dot{\mathbf{X}}_{\tau,1-\mathbf{m}}\,\Big|\,\mathbf{X}_{\tau,1-\mathbf{m}}=\mathbf{y},\Xi(\mathbf{u},\mathbf{w},\mathbf{m})\Bigr].$ (27)
We denote the density of random variable $\mathbf{X}_{\tau,1-\mathbf{m}}|\Xi(\mathbf{u},\mathbf{w},\mathbf{m})$ as $p_{\tau}(\mathbf{y}|\mathbf{u},\mathbf{w},\mathbf{m}).$
@@
We can see that due to the independence $\mathbf{X}_{1}\perp\!\!\!\!\perp(\mathbf{X}_{0},\mathbf{M}_{0})$,
$\mathbf{X}_{0,1-\mathbf{m}}|\mathbf{X}_{0,\mathbf{m}},\mathbf{X}_{1,\mathbf{m}},\mathbf{M}_{0}=\mathbf{X}_{0,1-\mathbf{m}}|\mathbf{X}_{0,\mathbf{m}},\mathbf{M}_{0},$ (28)
@@
$\mathbf{X}_{1,1-\mathbf{m}}|\mathbf{X}_{0,\mathbf{m}},\mathbf{X}_{1,\mathbf{m}},\mathbf{M}_{0}=\mathbf{X}_{1,1-\mathbf{m}}|\mathbf{X}_{1,\mathbf{m}},$ (29)
@@
The density function $p_{\tau}(\mathbf{y}|\mathbf{u},\mathbf{w},\mathbf{m})$ satisfies the continuity equation
$\partial_{\tau}p_{\tau}(\mathbf{y}|\mathbf{u},\mathbf{w},\mathbf{m})+\nabla_{\mathbf{y}}\cdot\Big(p_{\tau}(\mathbf{y}|\mathbf{u},\mathbf{w},\mathbf{m})\,\mathbf{v}^{*}(\mathbf{y},\mathbf{u},\mathbf{w},\mathbf{m},\tau)\Big)=0$ (30)
@@
$\mathbb{E}\big[h(\mathbf{X}_{\tau,1-\mathbf{m}})|\Xi(\mathbf{u},\mathbf{w},\mathbf{m})\big]=\int_{\mathbb{R}^{d_{1-\mathbf{m}}}}h(\mathbf{y})\,p_{\tau}(\mathbf{y}|\mathbf{u},\mathbf{w},\mathbf{m})\,\mathrm{d}\mathbf{y}$ (31)
@@
$\partial_{\tau}\mathbb{E}\big[h(\mathbf{X}_{\tau,1-\mathbf{m}})|\Xi(\mathbf{u},\mathbf{w},\mathbf{m})\big]=\int h(\mathbf{y})\,\partial_{\tau}p_{\tau}(\mathbf{y}|\mathbf{u},\mathbf{w},\mathbf{m})\,\mathrm{d}\mathbf{y}.$ (32)
On the other hand, we can see that $\mathbf{X}_{\tau}$ is a deterministic function of $\mathbf{X}_{0}$ and $\mathbf{X}_{1}$, thus using the law of unconscious statistician:
$\partial_{\tau}\mathbb{E}\big[h(\mathbf{X}_{\tau,1-\mathbf{m}})|\Xi(\mathbf{u},\mathbf{w},\mathbf{m})\big]$ (33)
$=$ $\partial_{\tau}\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{1}}\big[h(\mathbf{X}_{\tau,1-\mathbf{m}})|\Xi(\mathbf{u},\mathbf{w},\mathbf{m})\big]$ (34)
$=$ $\mathbb{E}_{\mathbf{X}_{0},\mathbf{X}_{1}}\big[\partial_{\tau}h(\mathbf{X}_{\tau,1-\mathbf{m}})|\Xi(\mathbf{u},\mathbf{w},\mathbf{m})\big]$ (35)
$=$ $\mathbb{E}\big[\partial_{\tau}h(\mathbf{X}_{\tau,1-\mathbf{m}})|\Xi(\mathbf{u},\mathbf{w},\mathbf{m})\big].$ (36)
@@
$\partial_{\tau}\mathbb{E}\big[h(\mathbf{X}_{\tau,1-\mathbf{m}})|\Xi(\mathbf{u},\mathbf{w},\mathbf{m})\big]$ (37)
$=$ $\mathbb{E}\big[\partial_{\tau}h(\mathbf{X}_{\tau,1-\mathbf{m}})|\Xi(\mathbf{u},\mathbf{w},\mathbf{m})\big]$ (38)
@@
Now we equate the two representations ((32) and (43)) for $\partial_{\tau}\mathbb{E}[h(\mathbf{X}_{\tau,1-\mathbf{m}})|\Xi(\mathbf{u},\mathbf{w},\mathbf{m})]$:
@@
Since this equality holds for every smooth, compactly supported test function $h(\mathbf{y})$, the fundamental lemma of the calculus of variations implies that,
@@
We can see that with the initial condition
$\mathbf{Z}_{0}=\mathbf{X}_{0,1-\mathbf{m}}|\Xi(\mathbf{u},\mathbf{w},\mathbf{m}),$
Lemma 3 says
$\mathbf{Z}_{1}=\mathbf{X}_{1,1-\mathbf{m}}|\Xi(\mathbf{u},\mathbf{w},\mathbf{m}).$
Using Lemma 2, we simplify $\mathbf{Z}_{0}$ and $\mathbf{Z}_{1}$ as
$\mathbf{Z}_{0}=\mathbf{X}_{0,1-\mathbf{m}}|\left\{\mathbf{X}_{0,\mathbf{m}}=\mathbf{u},\mathbf{M}_{0}=\mathbf{m}\right\},~~~\mathbf{Z}_{1}=\mathbf{X}_{1,1-\mathbf{m}}|\left\{\mathbf{X}_{1,\mathbf{m}}=\mathbf{w}\right\}.$
Letting $\mathbf{u}=\mathbf{w}=\mathbf{x}_{\mathbf{m}}$, we obtain the desired results in Lemma 1.
Appendix E Proof of Proposition 2
@@
Since the sequential imputer (3) never changes the observed part $\mathbf{X}_{\mathbf{m}}^{(t)}$ nor the missing mask $\mathbf{M}$, we have
@@
where the marginal density $p(\mathbf{x}_{\mathbf{m}},\mathbf{m})$ does not depend on $\mathbf{g}$. Now we show that the KL divergence only depends on $\mathbf{g}$ through the KL divergence $\mathrm{D}[p_{\mathbf{g}}(\cdot|\mathbf{x}_{1-\mathbf{m}},\mathbf{m})|q(\cdot|\mathbf{x}_{1-\mathbf{m}})]$.
$\mathrm{D}\left[\mathbb{P}_{\mathbf{X}^{(t)}(\mathbf{g}),\mathbf{M}}\|\mathbb{P}_{\mathbf{X}^{(t-1)}}\otimes\mathbb{P}_{\mathbf{M}}\right]$ (47)
$=$ $\mathbb{E}_{(\mathbf{x},\mathbf{m})\sim(\mathbf{X}_{\mathbf{g}}^{(t)},\mathbf{M})}\left[\log\frac{p_{\mathbf{g}}(\mathbf{x},\mathbf{m})}{q(\mathbf{x})p(\mathbf{m})}\right]$ (48)
$=$ $\mathbb{E}_{(\mathbf{x},\mathbf{m})\sim(\mathbf{X}_{\mathbf{g}}^{(t)},\mathbf{M})}\left[\log\frac{p_{\mathbf{g}}(\mathbf{x}_{1-\mathbf{m}},\mathbf{x}_{\mathbf{m}},\mathbf{m})}{q(\mathbf{x}_{1-\mathbf{m}},\mathbf{x}_{\mathbf{m}})p(\mathbf{m})}\right]$ (49)
$=$ $\mathbb{E}_{(\mathbf{x},\mathbf{m})\sim(\mathbf{X}_{\mathbf{g}}^{(t)},\mathbf{M})}\left[\log\frac{p_{\mathbf{g}}(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}},\mathbf{m})}{q(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}})}\right]+\mathbb{E}_{(\mathbf{x},\mathbf{m})\sim(\mathbf{X}_{\mathbf{g}}^{(t)},\mathbf{M})}\left[\log\frac{p(\mathbf{x}_{\mathbf{m}},\mathbf{m})}{q(\mathbf{x}_{\mathbf{m}})p(\mathbf{m})}\right]$ (50)
$=$ $\sum_{\mathbf{m}\in\{0,1\}^{d}}p(\mathbf{m})\mathbb{E}_{\mathbf{x}\sim\mathbf{X}_{\mathbf{g}}^{(t)}|\mathbf{M}=\mathbf{m}}\left[\log\frac{p_{\mathbf{g}}(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}},\mathbf{m})}{q(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}})}\right]+\mathrm{const.}$ (51)
$=$ $\sum_{\mathbf{m}\in\{0,1\}^{d}}p(\mathbf{m})\underbrace{\int p(\mathbf{x}_{\mathbf{m}}|\mathbf{m})\mathrm{D}[p_{\mathbf{g}}(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}},\mathbf{m})\|q(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}})]\mathrm{d}\mathbf{x}_{\mathbf{m}}}_{A_{\mathbf{g}}}+\mathrm{const.},$ (52)
where $\mathrm{const.}$ does not depend on $\mathbf{g}$. The constant in (51) is due to the factorization (46):
$\mathbb{E}_{(\mathbf{x},\mathbf{m})\sim(\mathbf{X}_{\mathbf{g}}^{(t)},\mathbf{M})}\left[\log\frac{p(\mathbf{x}_{\mathbf{m}},\mathbf{m})}{q(\mathbf{x}_{\mathbf{m}})p(\mathbf{m})}\right]=\mathbb{E}_{p(\mathbf{x}_{\mathbf{m}},\mathbf{m})}\mathbb{E}_{p_{\mathbf{g}}(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}},\mathbf{m})}\left[\log\frac{p(\mathbf{x}_{\mathbf{m}},\mathbf{m})}{q(\mathbf{x}_{\mathbf{m}})p(\mathbf{m})}\right]=\mathrm{const.}$ (53)
Assuming $p(\mathbf{x}_{\mathbf{m}}|\mathbf{m})$ is positive everywhere, $A_{\mathbf{g}}$ is minimized if and only if
$\mathrm{D}[p_{\mathbf{g}}(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}},\mathbf{m})|q(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}})]=0.$ (54)
The KL divergence $\mathrm{D}[p_{\mathbf{g}}(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}},\mathbf{m})\|q(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}})]$ is zero if and only if
$p_{\mathbf{g}}(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}},\mathbf{m})=q(\mathbf{x}_{1-\mathbf{m}}|\mathbf{x}_{\mathbf{m}}).$ (55)